# MSCA-Net: Multi-Scale Context Aggregation Network for Infrared Small Target Detection

### v1

@@ IV-E Ablation study
In this section, we validate the effectiveness of the MSCA-Net model through baseline experiments. Table II shows the ablation study results for the proposed MSCA-Net model. We progressively added the MSEDA, PCBAM, and CA modules to the baseline model. The experiment was conducted on the NUAA-SIRST dataset.
TABLE II: Ablation results of MSCA-Net
MSEDA PCBAM CAB mIoU(%) nIoU(%) $P_{d}$(%) $F_{a}$($10^{-6}$)
@@
The experimental results show that the MSEDA module enhances the cross-level representation ability of small targets through a multi-scale feature fusion strategy, whereas the PCBAM module effectively focuses on key regions via a channel-space-location attention collaborative mechanism. The CAB module significantly improves the discriminative expression of features through global channel interaction. When these three components work together, they form a progressive optimization system involving scale adaptation, spatial localization, and channel enhancement. This synergy generates collaborative effects across three critical stages, i.e., feature extraction, attention selection, and feature enhancement, ultimately constructing a multi-level feature expression network with strong discriminative power for infrared small targets.
(1) Impact of the MSEDA module: The MSEDA module is designed for multi-scale feature extraction, thereby enhancing the overall feature representation of small targets. To validate its effectiveness, we compared the MSEDA module with two other multi-scale modules, SCTB and ASSA. The experimental results, shown in Table III, demonstrate that MSEDA outperforms both SCTB and the ASSA in improving the feature representation of small targets, leading to a significant increase in detection accuracy.
TABLE III: Comparative experiment of the MSEDA module in MSCA-Net
@@
ASSA [43] 76.07 76.19 96.93 22.71
MSEDA 76.71 78.84 95.44 24.63
(2) Impact of the PCBAM module: To validate its effectiveness, we compared it with two other attention modules, CFN and CBAM. The experimental results, presented in Table IV, indicate that the PCBAM module outperforms the other two modules in constructing contextual information and capturing the local feature space, thereby effectively improving the model’s detection accuracy.
TABLE IV: Comparative experiment of the PCBAM module in MSCA-Net
@@
CFN [35] 76.98 79.03 95.88 23.43
CBAM [41] 77.02 78.88 96.03 23.11
PCBAM 77.64 79.51 95.06 20.92
(3) Impact of the CAB module: To verify its effectiveness, we compared it with two other feature expression modules, the CCA and the FAM. The experimental results, shown in Table V, demonstrate that CAB outperforms the other two modules in aggregating information across different channels, enhancing the expression of critical features, and suppressing the interference of redundant features, thereby effectively improving the segmentation performance of the model.
TABLE V: Comparative experiments of CAB modules in MSCA-Net
@@ IV-F Visual results
On the NUAA-SIRST, NUDT-SIRST, and IRSTD-1K datasets, qualitative comparative analysis of seven mainstream infrared small target segmentation algorithms demonstrates that the method proposed in this study shows significant advantages under different clutter backgrounds and noise interference. This is attributed to the synergistic effect of our multi-scale feature extraction and position-channel-space attention mechanism, which effectively enhances the edge response features of small targets.
As shown in Figures 6 and 7, we selected six images with complex backgrounds and considerable interference for testing. The blue boxes indicate correct detections, the yellow boxes indicate false detections, and the red boxes indicate missed detections. The experimental results indicate that, compared with other algorithms, our method successfully detects every target, whereas other methods result in some missed and false detections. For example, as shown in Figure 6 (5), all methods, except for our method and DNA-Net, produced false alarms. This issue can be attributed to these methods relying solely on local contrast information and lacking the ability to model long-range dependencies in the image, making them prone to false positives in complex backgrounds. As shown in Figure 6 (6), our method successfully segmented and localized the target, whereas other deep learning methods failed to detect it. Figure 7 presents a 3D visualization of the saliency maps generated by different methods on the six test images.
Figure 6: Visual results were obtained by different IRSTD methods on the NUAA-SIRST, NUDT-SIRST, and IRSTD-1K datasets. The circles in blue, red, and yellow represent correctly detected targets, missed detections, and false alarms, respectively. (a) Input. (b) ACM. (c) ALCNet. (d)RDIAN. (e) DNA-Net. (f) ISTDU-Net. (g) UIU-Net. (h) MSCA-Net. (i) GT.
@@
V Summary and conclusions
This study proposes a multi-scale perception-based infrared small target detection algorithm, MSCA-Net, which is built upon the U-Net architecture and employs a triple-coordinated optimization mechanism to overcome performance bottlenecks. Specifically, in the encoder stage, a multiscale enhanced dual attention mechanism (MSEDA) is introduced, combined with a cascaded dilated convolution module to adaptively aggregate multi-scale information, thereby enhancing the model’s feature extraction capability. Moreover, a position-aware cross-layer attention block (PCBAM) is incorporated to refine target features across the spatial, channel, and positional dimensions, facilitating deep interactions between global and local information. In the decoder stage, a Channel Aggregation Block (CAB) is further integrated to optimize the allocation of input feature channels, enabling the efficient transmission of critical features and improving the model’s robustness in complex backgrounds. The experimental results demonstrate that MSCA-Net achieves outstanding detection performance on three benchmark datasets—NUAA-SIRST, NUDT-SIRST, and IRSTD-1K—where quantitative evaluations strongly validate its effectiveness and superiority.

### v2

@@ IV-E Ablation study
In this part, we validate the effectiveness of the MSCA-Net model through U-Net baseline experiments. Table III shows the ablation study results for the proposed MSCA-Net model. We progressively added the RBs (residual blocks), the DS (Deep Supervision), the MSEDA, PCBAM, and CA modules to the baseline model. The experiment was conducted on the NUAA-SIRST dataset.
TABLE III: Ablation results of MSCA-Net
U-Net +RBs +DS MSEDA PCBAM CAB mIoU(%) nIoU(%) $P_{d}$(%) $F_{a}$($10^{-6}$)
@@
The experimental results show that each module brings noticeable improvements. It is worth emphasizing that the incorporation of MSEDA, PCBAM, and CAB modules yields modest improvements in mIoU, nIoU, and $P_{d}$, with increases ranging from 0.40% to 1.29%. Meanwhile, the $F_{a}$ metric shows significant reductions of 10.1%, 12.9%, and 19.7%, respectively. These results further demonstrate the effectiveness and robustness of the proposed modules in enhancing detection accuracy and suppressing false alarms.To further analyze their contributions, we examine the roles of the individual components in detail.
The MSEDA module enhances the cross-level representation ability of small targets through a multi-scale feature fusion strategy, whereas the PCBAM module effectively focuses on key regions via a channel-space-location attention collaborative mechanism. CAB strengthens the integration of essential feature representations by emphasizing key channels through adaptive weighting. When these three components work together, they form a progressive optimization system involving scale adaptation, spatial localization, and channel enhancement. This synergy generates collaborative effects across three critical stages, i.e., feature extraction, attention selection, and feature enhancement, ultimately constructing a multi-level feature expression network with strong discriminative power for infrared small targets.
To further evaluate the effectiveness of the MSEDA, PCBAM, and CAB modules, comparative experiments are conducted against several SOTA counterparts for each module individually.
(1) Impact of the MSEDA module: The MSEDA component is developed to facilitate the extraction of features across multiple scales, thereby enhancing the model’s ability to capture both coarse and fine-grained information. To confirm its functionality, we compared the MSEDA module with two other multi-scale modules, SCTB and ASSA. The experimental results, presented in Table IV, demonstrate that MSEDA outperforms both SCTB and the ASSA in improving the feature representation of small targets, leading to a significant increase in detection accuracy.
TABLE IV: Comparative experiment of the MSEDA module in MSCA-Net
@@
ASSA [44] 76.07 76.19 95.03 24.71
MSEDA 77.31 78.94 95.44 24.03
(2) Impact of the PCBAM module: To validate its effectiveness, we compared it with two other attention modules, CFN and CBAM. The empirical findings, presented in Table V, indicate that the PCBAM module outperforms the other two modules in constructing contextual dependencies and capturing the local feature representations, thereby effectively enhancing the detection precision of the model.
TABLE V: Comparative experiment of the PCBAM module in MSCA-Net
@@
CFN [35] 76.98 79.03 95.08 23.43
CBAM [42] 77.02 78.88 94.93 23.11
PCBAM 77.64 79.51 95.93 20.92
(3) Impact of the CAB module. To verify its effectiveness, we compared it with two other feature fusion modules, the CCA and the FAM. The experimental results, shown in Table VI, demonstrate that CAB outperforms the other two modules in integrating essential representations by adaptively emphasizing key channels, strengthening informative content, and mitigating redundant interference, thus improving the model’s segmentation precision.
TABLE VI: Comparative experiments of CAB modules in MSCA-Net
@@
Figure 6: Visualization of the saliency maps were obtained by various IRSTD methods. The circles in blue, red, and yellow represent correctly detected targets, missed detections, and false alarms, respectively. (a) Input. (b) ACM. (c) ALCNet. (d)RDIAN. (e) DNA-Net. (f) ISTDU-Net. (g) UIU-Net. (h) MSCA-Net. (i) GT.
@@ IV-F Visual results
Using three publicly available IRSTD datasets, we qualitatively compare seven representative IRSTD algorithms. The proposed method demonstrates marked improvements in reducing missed and false detections under complex scenes, including targets partially obscured by haze in the sky, embedded within bright building structures, or nearly merged with dark backgrounds ect. These advantages stem from the combined effect of multi-scale feature extraction and a position–channel–spatial attention mechanism, which jointly enhances edge response and suppresses background interference. As a result, our method achieves clearer target delineation and significantly fewer false detections in diverse and cluttered environments.
As shown in Figures 6 and 7, we selected six images with complex backgrounds and considerable interference for testing. The blue boxes represent true positives, the yellow boxes denote false positives, and the red boxes highlight missed detections. The experimental results indicate that, compared with other algorithms, our method successfully detects every target, whereas other methods result in some missed and false detections. For instance, as shown in Figure 6 (5), all methods, except for our method and DNA-Net, produced false alarms. This issue can be attributed to these methods relying solely on local contrast information and lacking the ability to model long-range dependencies in the image, making them prone to false positives in complex backgrounds. As shown in Figure 6 (6), our method successfully segmented and localized the target, whereas other deep learning methods failed to detect it. Figure 7 displays a 3D representation of the saliency maps produced by various methods across the six test images. This is consistent with the visual results shown in Figure 6.
@@
V Conclusion
This study proposes a multi-scale perception-based infrared small target detection algorithm, MSCA-Net, which is based on the U-Net architecture and employs a triple-coordinated optimization mechanism to overcome performance bottlenecks. Specifically, in the encoder stage, a Multiscale Enhanced Dilated Attention is introduced, combined with a cascaded dilated convolution module to adaptively aggregate multi-scale information, thus strengthening the model’s ability to extract discriminative features. Moreover, a Positional Convolutional Block Attention is added to strengthen the feature response of key areas through the space-channel-position triple attention mechanism, promoting the coordinated representation optimization of local details and global semantics. In the decoder stage, a Channel Aggregation Feature Fusion Block is further incorporated to consolidate informative representations by adaptively highlighting pivotal channels. This facilitates effective feature propagation and enhances the model’s robustness under complex. The experimental outcomes reveal that MSCA-Net delivers remarkable detection capabilities across three benchmark datasets—NUAA-SIRST, NUDT-SIRST, and IRSTD-1K—where quantitative analyses robustly affirm its efficacy and advantages.